HUBBLE SPACE TELESCOPE DAILY REPORT #1167 PERIOD COVERED: 0700 EST Daily Status Report as of 175/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 5362 (A Time Resolved High Resolution UV Study of the Wind Outflow) The Faint Object Spectrograph (Blue Side) was used to make observations of the cataclysmic binary star V795-Herculis which is known to make spectrographic changes in the order of a few minutes and other longer periods. The onboard target acquisition was successful. The Science observations remain on Science Tape Recorder #3, and will be analyzed later. No anomalies were noted. 1.2 Completed WF/PC-2 5369 (HST Medium Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 5653 (Cycle 4 DD HST Observations of Comet Shoemaker-Levy (1993E) for June) The WF/PC-2 was used to make observations of the comet Shoemaker-Levy before its impact with the planet Jupiter to help refine the size estimates for the various components in the comet's train. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed FOC/96 5440 (Spectra Image of the Jovian Aurora) The Faint Object Camera (F/96) was used to obtain spectra of Jupiter's aurorae. Two observations were executed as scheduled, and the observations remain on Science Tape Recorder #3. They will be analyzed later. No anomalies were noted. 1.5 Completed WF/PC-2 5570 (WF/PC-2 Cycle 4 Calibration: Earthflats, Large Set) The WF/PC-2 was used to make a set of Earth flat calibration images using different filters in order to construct high quality flat fields of the OTA illumination pattern. The observations were executed as scheduled, and no anomalies have been noted. 1.6 Completed FOS/BL 5579 (Cycle 4 Test: To Determine FOS Polarimetry Mode) The Faint Object Spectrograph (Blue Side) was used to make calibration observations of its polarization capabilities by using standard star BD+28D4211. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled :9 Successful :9 2.3 Significant HSTARs: None 2.4 Operations Notes: During the recent operational day, there were no reported anomalies in any HST onboard subsystem. 3.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-